calculation. b. At what age can she retire if she puts 250/month into the account every month beginning one month from today? Please show your calculation. Cody estimates that he can invest with a return of 6% compounded annually, which of the following is the best investment opportunity? Please show your calculation. 2. a) to receive $100,000 today b) to receive $400,000 at the end of the 15th year c) to receive $18,000 at the end of each year for 10 years d) to receive $75,000 at the end of 5th year and $100,000 at the end of 10th year e) to receive $75,000 at the end of 5th year and $175,000 at the end of 15th year 1) Use the items below to prepare a personal balance sheet for Jim Gordon (use the worksheet Topic3 BalanceSheet posted on D2L). Hint: not all information is relevant. 3.
Answer #1

Question 2.

Solution: For making determining the best investment opportunity we have calculate the present value of each alternative:

Present value of alternative (a) = $100,000 (given)

Present value of alt. (b) = $400,000 / (1.06)15 = 166,906.02

Present value of alt. (c) = 18,000 [1-(1.06)-10] / 0.06 = 132,481.57

Present value of alt. (d) = $75,000/ (1.06)5 + $100,000/(1.06)10 = 111,883.84

Present value of alt. (e) = $75,000 /(1.06)5+ $175,000 / (1.06)15 = $129,065.74

Since, Present value of the alternative (b). is higher than others thus, it is a best investment opportunity.